OpenStudy (anonymous):

A=1/2h(b+B) A=91 h=7 b=12

OpenStudy (anonymous):

solve for B i assume. 3 words. PLUG AND CHUGGGGGG

OpenStudy (anonymous):

How do I solve for B

OpenStudy (anonymous):

you would do 1/2 times 7 then distribute that answer to the b and the B

OpenStudy (anonymous):

How do you multiply 1/2 times 7 that will equal 7/2

OpenStudy (anonymous):

make it into a decimal and do .5 times 7

OpenStudy (anonymous):

can you show me in steps how it would look

OpenStudy (anonymous):

91=1/2(7)(12+B) 91=35(12+B) 91=420(B)